x-y= 6
3x+5y=30

answer:( 7 1/2, 1 1/2)
need a full break down on how my text book came up with these solutions

Answer by checkley77(12569)

x-y=6 multiply by 5 & add.
3x+5y=30
5x-5y=30
-------------------
8x=60
x=60/8
x=7.5 ans.
7.5-y=6
-y=6-7.5
-y=-7.5+6
-y=-1.5
y=1.5 ans.